Circular No. 6345

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) M. Womack, Pennsylvania State University; M. C. Festou, Observa-

toire Midi Pyrenées, Toulouse; and S. A. Stern, Southwest Research Institute, report observations of comet C/1996 B2 with the National Radio Astronomy Observatory’s 12-m telescope at Kitt Peak: “The following lines were detected during Mar. 16 and 17: H_{2}CO 3(12)-2(11) at 226 GHz, HCN 3-2 at 266 GHz, CO 2-1 at 230 GHz and CH_{3}OH at 145 GHz. Based on the detected brightness temperatures and an assumed rotational temperature of 50 K we estimate production rates of Q(H_{2}CO) = 2 × 10^{26}/s, Q(CO) = 3 × 10^{28}/s, Q(CH_{3}OH) = 5 × 10^{26}/s and Q(HCN) = 3 × 10^{26}/s. This Q(HCN) is close to that found for CN by R. Millis and D. Schleicher on Mar. 19 (IAUC 6344), suggesting a direct link between the two species. However, the CO and H_{2}CO data indicate that H_{2}CO is not the main parent of the observed CO. Interpolation between the present observations of CO and those reported on IAUC 6335 (and on the assumption that the fluorescence efficiencies of the millimeter lines have not significantly changed) indicates a very rapid change in the production rates, presumably near an r^{−5} law, during the last two weeks. If the CO/H_{2}O ratio is 20 percent, as suggested on IAUC 6335, the water production is in excess of 10^{29}/s.”

Improved orbital elements and a detailed new ephemeris are given on MPEC 1996-F03. The following elements are from the same solution but for a more current osculation epoch:

Epoch = 1996 Mar. 18.0 TT

T = 1996 May 1.3966 TT  ω = 130.^{o}2016
e = 0.999705  Ω = 188.0432 }2000.0
q = 0.230085 AU  i = 124.9106

The earlier indication (IAUC 6329) that the comet is not “new” is confirmed, the last perihelion passage having occurred some 9000 years ago. Corrections to the IAUC 6330 ephemeris: Mar. 22, Δα = −0^{m}.01, Δδ = +3′.5; 23, −0^{m}.02, +5′.2; 24, −0^{m}.04, +7′.2; 25, −0^{m}.09, +9′.0; 26, −0^{m}.4, +9′.0; 27, −7^{m}, +3′.3; 28, −0^{m}.7, −5′.8; 29, −0^{m}.19, −4′.7; 30, −0^{m}.10, −3′.6; 31, −0^{m}.06, −2′.8; Apr. 1, −0^{m}.04, −2′.4; 2, −0^{m}.04, −2′.0; then to the IAUC 6311 ephemeris: Apr. 2, −0^{m}.08, −4′.4; 7, −0^{m}.05, −2′.3; 12, −0^{m}.04, −1′.7; 17, −0^{m}.05, −1′.4; 22, −0^{m}.06, −1′.4; 27, −0^{m}.08, −1′.7.

Naked-eye magnitude estimates: Mar. 20.07 UT, 1.9 (B. H. Granslo, Fjellhamar, Norway); 20.09, 2.3 (K. Černis, Lavariskes, Lithuania); 20.58, 2.5 (S. O’Meara, Volcano, HI); 20.98, 2.1 (P. Pravec, Ondřejov, Czech Republic); 21.16, 2.2 (B. G. Marsden, Lexington, MA); 21.31, 1.6 (J. V. Scotti, Tucson, AZ); 21.48, 1.3 (C. S. Morris, near Gorman, CA).

1996 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden